OpenStudy (anonymous):

Solve for u. u-4/5=6 3/4

OpenStudy (anonymous):

\[u-\frac{ 4 }{ 5}=6\frac{ 3 }{ 4 }\] is this what you mean?

OpenStudy (abb0t):

add \(\frac{4}{5}\) to both sides.

OpenStudy (anonymous):

Yes my bad

OpenStudy (anonymous):

what do you mean add 4/5 to both sides?

OpenStudy (abb0t):

Add!

OpenStudy (abb0t):

\(u - \frac{4}{5} + \frac{4}{5} = 6\frac{3}{4} + \frac{4}{5}\)

OpenStudy (abb0t):

change your mixed fraction into improper fraction.

OpenStudy (anonymous):